# skipabove=1em,skipbelow=1em

### v3

@@ 4.2.4. Safe Eating Practices
Safe eating practices aim to ensure that the ingestion process does not pose additional health risks. Current interventions in this area primarily target two aspects: hygiene behaviors and food temperature control. Hygiene-related interventions focus on promoting safe practices during eating to prevent foodborne infections (Wang and Zhang, 2021). Other interventions address food temperature control to prevent oral injuries, such as burns (Singhal et al., 2024).
5. Structures and Design Space for Behavioral Sensing and Intervention Paradigms
This chapter analyzes the structural and modality-level characteristics of ingestion health paradigms, as outlined in Chapter 3.2, with a focus on two essential stages of the behavioral loop: sensing and feedback/intervention. While reasoning processes typically mediate between perception and action, they are often embedded within system pipelines, reflected through how sensing data is interpreted and applied. Therefore, our analysis centers on the sensing and feedback/intervention components, where modality design plays a direct and observable role.
Inspired by Jaimes and Sebe (2007) and Jansen et al. (2022), we adopt modality as the classification standard for both the sensing and intervention components, encompassing interactive devices and perceptual computing methods. Specifically, Jaimes and Sebe (2007) define modality as “mode of communication according to human senses and computer input devices activated by humans or measuring human qualities,” emphasizing that modality serves as a communication channel designed around human perception and computational input-output mechanisms. Furthermore, they introduce a human-based perspective on modality, identifying core modalities corresponding to human senses, including vision, touch, hearing, smell, and taste, and categorizing more complex examples within these sensory categories. Building on this foundation, Jansen et al. (2022) expand the concept of human-based modality by introducing two additional categories—cerebral and cardiac modalities—to accommodate advances in interactive device technologies.
Following insights from Jaimes and Sebe (2007), we recognize that, in the sensing process, certain modalities do not directly correspond to individual human sensory channels but involve interactions mediated through external devices or environmental signals. Thus, we introduce the category of environment-based modalities, capturing input and feedback mechanisms targeting environmental objects independently of human senses. This extension provides a more comprehensive classification for sensing and intervention designs in ingestion health systems, covering modalities arising from device manipulation, environmental interactions, or indirect physiological monitoring.
We adapt and refine this modality classification to suit ingestion health interventions, distinctly addressing both sensing and feedback/intervention stages. The primary classification human-based modality contains visual, auditory, haptic, gustatory, olfactory, motion, and physiological, and environment-based modality contains Visual, Acoustic, Force, Spectral, and Electrical. Almost each modality further includes submodalities considering focus and implications different from sensing and feedback/intervention. Accordingly, slight structural and descriptive variations in modality classifications are introduced to align precisely with each phase of the behavioral closed-loop, as shown in following subsections.
@@ 5.1. Sensing
This subsection focuses on the sensing modalities within overall ingestion health intervention paradigms, where behavioral data are primarily acquired through various sensor-based or perceptual computing-driven modalities. The detailed structure and descriptions of each modality, along with its associated submodalities in the sensing process, are summarized in Table 1.
Notably, gustatory and olfactory modalities are absent in the sensing stage across existing systems. We believe that sensing gustatory or olfactory information through perceptual or sensor-based means remains technically challenging, and these modalities are mainly employed in the feedback/intervention phase rather than for direct behavioral sensing.
The following sections present the identified sensing modalities, categorized according to their classifications, with representative approaches within ingestion health intervention paradigms.
Modality Submodality Description Paper count
[…]
@@ 5.1.1. Human-based
Visual: Visual sensing has been widely adopted in ingestion-related systems to extract behavioral and attentional cues through camera-based observation. We identified three common submodalities within visual sensing: gaze, facial features, and body movement.
@@
Gaze-based sensing tracks eye movements and fixation points to assess users’ attention during food-related interactions. For example, Karkar et al. (2018) integrated an eye tracker into a VR system to analyze children’s visual attention toward nutrition-themed stimuli. Webcam-based gaze estimation has also been used to infer attentional disengagement during meals, which can inform timely interventions (Khot et al., 2020).
Facial-expression-based sensing is primarily used to detect emotional or behavioral responses and confirm user presence. A webcam was employed to extract facial expressions and assess affective states during meals (Mayumi et al., 2023). Moreover, Kim and Bae (2018a) used lip and jaw motion analysis to model bite rhythm. Face detection has also supported social presence recognition in multi-user contexts, helping to determine whether a child is present (Triantafyllidis et al., 2019) or enabling socially responsive camera-based interaction (Robinson et al., 2020).
Body-based visual sensing targets ingestion-related gestures and postures, such as hand movements and upper-body orientation. Xie et al. (2023) tracked caregivers’ feeding gestures to estimate feeding frequency and potential satiety cues in children. Skeleton tracking with depth cameras has been used to monitor upper-body posture during meals (Mayumi et al., 2023). Additionally, egocentric video data were collected for post-hoc analysis of mealtime behaviors and social dynamics (Zhang and Parker, 2020).
@@
Auditory: Auditory sensing in ingestion health systems typically involves capturing either verbal input or eating-related sounds, which can be classified into two main submodalities: speech and non-verbal sound, covering existing approaches.
@@
Speech-based auditory sensing is commonly used for semantic data acquisition or enabling conversational interfaces. Spoken input has been leveraged to extract structured dietary information using linguistic pattern recognition (Prasetyo et al., 2020). Systems such as those proposed by Mayumi et al. (2023) and Robinson et al. (2020) receive users’ queries for food-related facts or recommendations and extract key information from speech signals. Speech has also been employed as an input modality for logging intake events or interacting with nutrition tracking systems (Luo et al., 2021; Arefeen et al., 2022). Parra et al. (2023) used Wizard-of-Oz setups and processed users’ responses for further system-guided prompting.
Sound-based sensing focuses on detecting physiological cues, such as chewing or swallowing, using microphones. Some researchers have placed microphones near the throat (Kalantarian et al., 2014) or collar (Chen and Yen, 2022) to capture oral sound signals during ingestion, which can be analyzed for bite detection (Bi et al., 2016), food classification (Kleinberger et al., 2023), or intake rhythm estimation. Microphones have also been implemented in earpiece form factors to detect rhythmic mastication events, such as chewing sensors (Maramis et al., 2014) and bone conduction microphones (Kamachi et al., 2023), which capture sound with reduced ambient interference.
Additionally, although categorized separately above, some microphones are designed for specific contexts, such as capturing chewing in denture wearers (Koizumi et al., 2011) or jointly recording speech and chewing sounds as contextual data (Zhang and Parker, 2020; Kamachi et al., 2023).
Haptic: Haptic sensing aims to detect physical interactions or internal mechanical activity through pressure, force, or contact-based sensors. A common form is weight sensing, which tracks changes in body weight as a long-term reflection of dietary intake (Ludwig et al., 2011; Bastida et al., 2023; Nedungadi et al., 2018).
@@
A substantial body of work has focused on haptic sensing using piezoelectric or force sensors. Neck-mounted piezoelectric sensors have been embedded into wearable necklaces to capture throat vibrations associated with swallowing (Alshurafa et al., 2015; Kim et al., 2016b; Abisha and Rajalakshmy, 2017). Similar techniques have been used to detect mechanical displacement in the temporalis muscle to infer chewing activity (Farooq et al., 2017). Other studies have explored intraoral pressure detection, such as placing sensors on palatal retainers to measure tongue force (Biyani et al., 2019), or using external force sensors to estimate masticatory force (Zheng et al., 2025). Additionally, pressure-sensitive buttons or interfaces have been applied to allow users to log ingestion-related events (Fortmann et al., 2014; Tang et al., 2015; Cai et al., 2024; Robinson et al., 2020).
@@
Motion: Motion sensing captures gestures and subtle physical activity using inertial or proximity sensors.
@@
Most gesture-oriented systems employ inertial measurement units (IMUs) mounted on utensils (Wang et al., 2025; Khot et al., 2020; Zhang et al., 2019; Li and Bailenson, 2018; Kadomura et al., 2013a, 2014), wristbands (Kim et al., 2016a; Kim and Bae, 2018b; Ye et al., 2016; Turner-McGrievy et al., 2017; Mendi et al., 2013; Kumbhare et al., 2023; Gaikwad et al., 2024; Dragoni et al., 2017; Scisco et al., 2011), or containers (Tommy et al., 2017; Bobin et al., 2018; Chiu et al., 2009; Zhou et al., 2021) to detect hand-to-mouth movement, bite frequency, or drinking actions. IMUs have also been integrated into smart forks (Hermans et al., 2017), smart glasses (Bedri et al., 2022), and smartphones (Tirasirichai et al., 2018). Additionally, some platforms or frameworks use pedometers (Harous et al., 2017) and activity trackers (Alexander et al., 2017; Vazquez et al., 2015; Fuchs et al., 2019) to estimate ingestion or physical activity context. These approaches have made it easier for broader health-monitoring systems to achieve dietary reasoning support (Bastida et al., 2023; Mishra et al., 2024; Maramis et al., 2014; Chukwu, 2011; Badawi et al., 2012a; Jung et al., 2007).
@@
Chewing and micro-head motion sensing focus on jaw, throat, or mouth-level interactions. Biyani et al. (2019) used a 3D accelerometer to track throat motion, and Hori et al. (2023) applied infrared distance sensing to capture subtle skin vibrations. Additionally, electrical conductivity sensors (Kadomura et al., 2013a, b) or static charge sensors (Han and Kang, 2017) were embedded in utensils or cups to infer mouth contact and drinking actions.
@@
Physiological: Physiological sensing provides insight into users’ internal states that influence or contextualize ingestion behavior, including appetite, metabolic status, stress, and cardiovascular conditions.
@@
Autonomic sensing typically involves EMG electrodes to detect muscle activity during chewing or wrist motion (Lee et al., 2019; Sugita et al., 2018; Kim et al., 2016b), as well as thermistors and respiratory sensors to track body temperature and breathing patterns (Park and Min, 2007; Jung et al., 2007). Emotional arousal or stress is commonly inferred through electrodermal activity (EDA) sensors (Carroll et al., 2013; Nedungadi et al., 2018), which may affect eating pace or impulsive behaviors.
@@
Cardiac sensing captures heart rate, blood pressure, and oxygen saturation using ECG (Badawi et al., 2012b), pulse wave (Varun et al., 2023), or optical PPG sensors (Park and Min, 2007; Ludwig et al., 2011; Jung et al., 2007; Huang et al., 2018; Nedungadi et al., 2018), which are frequently used to contextualize ingestive behavior or guide personalized intervention timing.
Others: Apart from the sensing modalities detailed above, there are several approaches that do not involve sensing or perceptual computing methods. These approaches rely entirely on user-reported data collected through manual input. Common input methods include the manual entry of dietary events or physical status, image selection or annotation, and questionnaire-style logging via mobile or chatbot interfaces. Although these systems may still provide feedback, they fall outside our sensing modality framework due to the lack of automated perceptual input or computational sensing logic. A detailed overview of this manual approach is summarized in Table 7, with approximately 20 publications adopting this method.
@@ 5.1.2. Environment-based
Visual: Environment-based visual sensing focuses on object-level perception of food and eating-related artifacts to support dietary monitoring and context acquisition. Most systems utilize food image analysis, applying computer vision to classify food types, estimate portion sizes, or infer nutritional content from captured images (Kanjalkar et al., 2024; Eigen et al., 2018; Cordeiro et al., 2015; YR et al., 2024; Epstein et al., 2016). These methods are commonly implemented through smartphone photography (Mishra et al., 2024; Waltner et al., 2015; Achilleos et al., 2017), utensil-mounted or refrigerator cameras (Nakamura et al., 2023; Kwon et al., 2016; Faltaous et al., 2021), or wearable vision modules (Fuchs et al., 2019).
@@
Some systems use food color (Kadomura et al., 2013a, 2014), utensil-based visual cues (Wang and Zhang, 2021), or visual distortion sensing (Akter et al., 2021; Chiu et al., 2009) to infer consumption context or eating correctness. Other approaches incorporate plate-level dynamics (Ganesh et al., 2014), receipt parsing (Mankoff et al., 2002), puzzle-based interaction (Cai et al., 2024), or packaging and interface-based food input (Arefeen et al., 2022; Hsiao et al., 2011) for supplemental inference and logging. Additionally, visual code recognition is widely used to access structured nutritional information through barcode scanning (Volkova et al., 2016; Eyles et al., 2017), QR code interpretation via mobile interfaces (Fernandez et al., 2013; Vazquez et al., 2015; Alloghani et al., 2016), and edible marker detection in AR-enhanced scenarios (Gutiérrez et al., 2018).
@@
Acoustic: Acoustic sensing often utilizes ultrasonic distance sensors to monitor liquid intake or food-related actions. Most related works involve embedding ultrasonic transducers into smart bottles or storage containers to measure fluid levels and infer hydration behavior (Poddar et al., 2024; Wijanarko et al., 2019; Ravindran et al., 2022; Varun et al., 2023; Tommy et al., 2017). These systems estimate drinking volume based on real-time fluid height or remaining capacity. Other systems apply similar sensing techniques to detect discrete ingestion-related events, such as the removal of a snack from a container using ultrasonic signal disruption (Liu et al., 2024), or monitoring food depletion in a storage cabinet based on internal distance measurements (Faltaous et al., 2021).
@@
Force: Force sensing typically involves the use of load cells or pressure sensors embedded in plates, coasters, or tables to measure changes in food or liquid weight over time. This approach enables passive monitoring of eating dynamics, such as intake amount, bite frequency, or consumption rate. Smart plates equipped with load cells (Mitchell et al., 2015; Kim et al., 2016c; Maramis et al., 2014; Ioakimidis et al., 2009; de Vries et al., 2020) or force-sensitive resistors (FSRs) (Jung et al., 2017) have been used to continuously track food depletion and infer eating pace or satiety.
Beyond intake tracking, similar sensing mechanisms have been extended to estimate nutrient intake across meals (Krishna et al., 2025; Zhao et al., 2021; de Vries et al., 2020; Jung et al., 2017), monitor liquid consumption via drink weight changes (Ritschel et al., 2018; Lessel et al., 2016), or support playful interactions through weight-sensitive input (Lo et al., 2007). Additionally, multi-plate setups have been applied to detect food preferences and portion control behaviors (Joi et al., 2016; Hermans et al., 2017).
@@
Spectral: Spectral sensing leverages temperature, proximity, and light-based sensors to capture physical or chemical properties of food, beverages, and environmental conditions. Common sensing components include infrared proximity detectors, temperature sensors, and light-spectrum analyzers. Specifically, infrared or optical sensors have been used to detect the presence of containers (Ritschel et al., 2018), estimate proximity to food (Bedri et al., 2022), or assess food temperature (Singhal et al., 2024). Temperature sensors embedded in bottles or utensils help monitor liquid warmth (Tommy et al., 2017; Wijanarko et al., 2019; Varun et al., 2023) or environmental conditions affecting food stability (Bastida et al., 2023; Badawi et al., 2012b; Kwon et al., 2016). In more advanced cases, Matsui et al. (2018) applied spectral absorption analysis to estimate the alcohol concentration in beverages and assess drinking speed.
@@
Electrical: Electrical sensing supports ingestion-related monitoring by detecting changes in conductivity, capacitance, or magnetic response, allowing for the inference of liquid consumption, food identity, or environmental conditions. This method enables passive, contact-based tracking of ingestion without requiring user instrumentation. Common sensor types include conductivity probes, humidity sensors, capacitive pads, and RFID/NFC modules.
@@
Specifically, conductivity and humidity sensors have been embedded into bottles or utensils to monitor water levels or detect drinking events (Bobin et al., 2018; Tommy et al., 2017; Kreutzer et al., 2015; Varun et al., 2023), while capacitive level sensing has been used to estimate water volume based on changes in dielectric properties (Kaner et al., 2018). Some systems integrate float switches or conductive tilt detection for recognizing bottle use (Pankajavalli et al., 2017; Ko et al., 2007). Ambient humidity is also used to characterize dining environments and inform food safety assessments (Bastida et al., 2023; Badawi et al., 2012b; Kwon et al., 2016).
@@
On the magnetic side, RFID and NFC technologies enable automatic identification of food items, storage states, or utensil presence (Vazquez et al., 2015; Huang et al., 2010; Jung et al., 2017), and advanced NFC designs even estimate consumed volume through passive milliliter tracking (Yıldız and Coşkun, 2019). Additionally, Han and Kang (2017) used metal contact detection to support real-time eating activity sensing.
5.2. Feedback and Intervention
As the final stage in the ingestion behavior loop, feedback and intervention mechanisms convey information back to users to support behavioral awareness, regulation, or change. In contrast to sensing, which focuses on data acquisition through sensors or perceptual computing, intervention emphasizes human interpretation and experience. Accordingly, the modalities at this stage are inherently human-based, as they operate by stimulating human sensory channels to deliver prompts, guidance, or other affective cues.

### v4

@@ 4.2.4. Safe Eating Practices
Safe eating practices are intended to minimize health risks during the ingestion process. Current interventions primarily target two aspects: hygiene behaviors and food temperature control. Hygiene-related interventions focus on promoting safe practices during eating to prevent foodborne infections (Wang and Zhang, 2021). Other interventions address food temperature control to prevent oral injuries, such as burns (Singhal et al., 2024).
5. Structures and Design Space from Behavioral Sensing to Intervention Paradigms
This section analyzes the modality-level characteristics of ingestion health systems, focusing on two essential stages of the behavioral loop: sensing and intervention. While reasoning processes mediate between perception and action, they are typically embedded within pipelines, reflected in how sensing data is interpreted and applied. Therefore, our analysis centers on the sensing and intervention components, where the design of modality plays a direct and observable role in shaping user experience and effectiveness of the system.
Prior research has established several perspectives on modality. Jaimes and Sebe (2007) define modality as ”modes of communication according to human senses and computer input devices activated by humans or measuring human qualities”, emphasizing the centrality of perception and input-output channels. Building on this view, Jansen et al. (2022) extend the concept by introducing cerebral and cardiac modalities to accommodate emerging interactive technologies. However, in ingestion-related systems, certain modalities do not correspond directly to individual human sensory channels but instead involve environmental interactions or indirect physiological monitoring. To address this, we introduce the category of environment-based modalities, which encompass input and feedback mechanisms targeting environmental objects independently of human senses. This extension enables a more comprehensive classification of sensing and intervention designs in ingestion health interventions.
Building on these foundations, we classify human-based modalities into visual, auditory, haptic, gustatory, olfactory, motion, and physiological, and environment-based modalities into Visual, Acoustic, Force, Spectral, and Electrical. Slight structural and descriptive variations in modality classifications are introduced below to align precisely with each phase of the behavioral closed-loop.
@@ 5.1. Sensing
This subsection focuses on sensing modalities within overall ingestion health intervention paradigms, where behavioral data are primarily acquired through various sensor-based or perceptual computing-driven modalities. The detailed structure and descriptions of each modality, along with its associated submodalities in the sensing process, are summarized in Table 5.
Notably, gustatory and olfactory modalities are absent in the sensing stage across existing systems. We believe that sensing gustatory or olfactory information through perceptual or sensor-based means remains technically challenging, and these modalities are mainly employed in the intervention phase rather than for direct behavioral sensing. The identified sensing modalities is categorized according to their classifications below. A summary of sensing modalities and corresponding publications included in this review is shown in Table 3.
@@ 5.1.1. Human-based
i. Visual: Visual sensing has been widely adopted in ingestion-related systems to extract behavioral and attentional cues through camera-based observation. We identified three common submodalities:
@@
Gaze sensing tracks eye movements and fixation points to assess users’ attention during food-related interactions. It has been used to measure attentional engagement with food stimuli across settings such as VR eye tracking (Karkar et al., 2018) or webcam-based monitoring (Khot et al., 2020).
Facial-expression sensing primarily support both emotional and behavioral inference in ingestion-related systems. It has been used to detect affective states (Mayumi et al., 2023), model chewing through lip and jaw motion (Kim and Bae, 2018a), and recognize user presence in multi-user contexts (Triantafyllidis et al., 2019; Robinson et al., 2020).
Body posture sensing interprets behaviors by capturing body and head orientation, alignment, and movement during ingestion. Posture-related signals were captured by tracking the feeding gestures of children for satiety signals (Xie et al., 2023), monitoring the upper-body orientation using skeleton models (Mayumi et al., 2023), and analyzing egocentric video for mealtime behavior (Zhang and Parker, 2020).
@@
ii. Auditory: Auditory sensing in ingestion health systems typically captures either verbal input or eating-related sounds, which can be classified into two main submodalities:
@@
Speech sensing commonly supports semantic data acquisition and conversational interfaces. Spoken input has been used to extract structured dietary data via linguistic pattern recognition (Prasetyo et al., 2020), facilitate food-related recommendations from spoken queries (Mayumi et al., 2023; Robinson et al., 2020), and log intake events with interaction (Luo et al., 2021; Arefeen et al., 2022), as well as to enable system-guided prompting via Wizard-of-Oz protocols (Parra et al., 2023).
Non-verbal sound sensing focuses on physiological cues like chewing or swallowing. Microphones have been placed near the throat (Kalantarian et al., 2014) or collar (Chen and Yen, 2022) to detect bites (Bi et al., 2016), classify food types (Kleinberger et al., 2023), or estimate intake rhythms. Earpiece-mounted sensors, including chewing microphones (Maramis et al., 2014) and bone conduction devices (Kamachi et al., 2023), help reduce ambient noise. Some systems further adapt to context, such as denture-based chewing detection (Koizumi et al., 2011) or setups capturing both speech and eating sounds (Zhang and Parker, 2020; Kamachi et al., 2023).
@@
iii. Haptic: Haptic sensing aims to detect physical interactions or internal mechanical activity through pressure, force, or contact-based sensors. A common application is weight sensing, which serves as a long-term indicator of dietary intake (Ludwig et al., 2011; Bastida et al., 2023; Nedungadi et al., 2018). Wearable piezoelectric and force sensors have been used to capture throat vibrations during swallowing (Alshurafa et al., 2015; Kim et al., 2016b; Abisha and Rajalakshmy, 2017) and temporalis muscle displacement for chewing estimation (Farooq et al., 2017). Other methods include intraoral pressure sensors embedded in palatal retainers to measure tongue force (Biyani et al., 2019), external sensors for estimating masticatory force (Zheng et al., 2025), and pressure-sensitive interfaces for logging ingestion events (Fortmann et al., 2014; Tang et al., 2015; Cai et al., 2024; Robinson et al., 2020).
@@
iv. Motion: Motion sensing captures gestures and subtle physical activity using inertial or proximity sensors.
@@
Most gesture-oriented systems employ inertial measurement units (IMUs) mounted on utensils (Wang et al., 2025; Khot et al., 2020; Zhang et al., 2019; Li and Bailenson, 2018; Kadomura et al., 2013a, 2014; Hermans et al., 2017), containers (Tommy et al., 2017; Bobin et al., 2018; Chiu et al., 2009; Zhou et al., 2021), wristbands (Kim et al., 2016a; Kim and Bae, 2018b; Ye et al., 2016; Turner-McGrievy et al., 2017; Mendi et al., 2013; Kumbhare et al., 2023; Gaikwad et al., 2024; Dragoni et al., 2017; Scisco et al., 2011), smart glasses (Bedri et al., 2022) and smartphones (Tirasirichai et al., 2018) to detect hand-to-mouth movement, bite frequency, or drinking actions. Some platforms also use pedometers (Harous et al., 2017) and activity trackers (Alexander et al., 2017; Vazquez et al., 2015; Fuchs et al., 2019) to estimate ingestion or physical activity context. These approaches have made it easier for broader health-monitoring systems to achieve dietary reasoning support.
@@
Chewing and micro-head motion sensing target jaw, throat, or mouth-level interactions. Biyani et al. (2019) tracked throat motion using a 3D accelerometer, while Hori et al. (2023) used infrared sensors to detect subtle skin vibrations. Sensors based on electrical conductivity (Kadomura et al., 2013a, b) or static charge (Han and Kang, 2017) have also been embedded in utensils and cups to infer mouth contact and drinking events.
@@
v. Physiological: Physiological sensing provides insight into users’ internal states that influence or contextualize ingestion behavior, including appetite, metabolic status, stress, and cardiovascular conditions.
@@
Autonomic sensing typically involves EMG electrodes to detect muscle activity during chewing or wrist motion (Lee et al., 2019; Sugita et al., 2018; Kim et al., 2016b), along with thermistors and respiratory sensors to track body temperature and breathing patterns (Park and Min, 2007; Jung et al., 2007). Emotional arousal or stress is inferred through electrodermal activity (EDA) sensors (Carroll et al., 2013; Nedungadi et al., 2018).
@@
Cardiac sensing captures heart rate, blood pressure, and oxygen saturation using ECG (Badawi et al., 2012b), pulse wave (Varun et al., 2023), or optical PPG sensors (Park and Min, 2007; Ludwig et al., 2011; Jung et al., 2007; Huang et al., 2018; Nedungadi et al., 2018) to support context-aware interpretation and intervention timing.
vi. Others: Without sensing or perceptual computing, other approaches rely entirely on manually collected self-reports. Common inputs include manual logging of dietary events or physical states, image selection or annotation, and questionnaire-style via mobile or chatbot interfaces. Though such systems can support reasoning and intervention, they fall outside our sensing stages due to the lack of automated perceptual input or computational sensing logic.
@@ 5.1.2. Environment-based
i. Visual: Environment-based visual sensing focuses on object-level perception of food and eating-related artifacts to support dietary monitoring and context acquisition. Most systems apply computer vision to classify food types, estimate portions sizes, or infer nutrition from images (Kanjalkar et al., 2024; Eigen et al., 2018; Cordeiro et al., 2015; YR et al., 2024; Epstein et al., 2016) ,via smartphone photography (Mishra et al., 2024; Waltner et al., 2015; Achilleos et al., 2017), utensil- or refrigerator-mounted cameras (Nakamura et al., 2023; Kwon et al., 2016; Faltaous et al., 2021), or wearable modules (Fuchs et al., 2019).
@@
Other methods use food color (Kadomura et al., 2013a, 2014), utensil-based cues (Wang and Zhang, 2021), visual distortion (Akter et al., 2021; Chiu et al., 2009), or plate dynamics (Ganesh et al., 2014) to infer consumption. Additional approaches include receipt parsing (Mankoff et al., 2002), puzzle-based interaction (Cai et al., 2024), packaging input (Arefeen et al., 2022; Hsiao et al., 2011), and visual code recognition such as barcode scanning (Volkova et al., 2016; Eyles et al., 2017), QR code interpretation (Fernandez et al., 2013; Vazquez et al., 2015; Alloghani et al., 2016), and edible marker detection in AR scenarios (Gutiérrez et al., 2018).
@@
ii. Acoustic: Acoustic sensing often utilizes ultrasonic distance sensors to monitor liquid intake or food-related actions. Most systems embed ultrasonic transducers in bottles or containers to track fluid levels and estimate drinking volume (Poddar et al., 2024; Wijanarko et al., 2019; Ravindran et al., 2022; Varun et al., 2023; Tommy et al., 2017). Others apply similar ultrasonic sensors to detect discrete events like snack removal (Liu et al., 2024), or monitoring food depletion in storage cabinets (Faltaous et al., 2021).
@@
iii. Force: Force sensing typically uses load cells or pressure sensors embedded to measure changes in food or liquid weight over time. Smart utensils equipped with load cells (Mitchell et al., 2015; Kim et al., 2016c; Maramis et al., 2014; Ioakimidis et al., 2009; de Vries et al., 2020; Lessel et al., 2016; Ritschel et al., 2018) or force-sensitive resistors (FSRs) (Jung et al., 2017) have been used to continuously track food or liquid depletion and infer ingestion pace or satiety. Building on ingestion tracking, systems aim to estimate nutrient intake (Krishna et al., 2025; Zhao et al., 2021; de Vries et al., 2020; Jung et al., 2017), detect food preferences and portion control (Joi et al., 2016; Hermans et al., 2017), and even gamify eating through weight-sensitive interactions (Lo et al., 2007).
@@
iv. Spectral: Spectral sensing commonly leverages infrared detectors, temperature sensors, and light-spectrum analyzers, to capture physical or chemical properties of food, beverages, and environmental conditions. Infrared or optical sensors have been used to detect the presence of containers (Ritschel et al., 2018), estimate proximity to food (Bedri et al., 2022), or assess food temperature (Singhal et al., 2024). Temperature sensors embedded in utensils help monitor liquid warmth (Tommy et al., 2017; Wijanarko et al., 2019; Varun et al., 2023) or environmental conditions affecting food stability (Bastida et al., 2023; Badawi et al., 2012b; Kwon et al., 2016). In addition, Matsui et al. (2018) used spectral absorption to estimate alcohol content and drinking rate.
@@
v. Electrical: Electrical sensing supports ingestion monitoring by detecting changes in conductivity, capacitance, or magnetic response. It enables passive contact-based tracking without user instrumentation. Common sensors include conductivity probes, humidity sensors, capacitive pads, and RFID/NFC modules.
@@
Conductivity and humidity sensors embedded in utensils can monitor water levels and drinking events (Bobin et al., 2018; Tommy et al., 2017; Kreutzer et al., 2015; Varun et al., 2023), while capacitive sensing can estimate volume via dielectric change (Kaner et al., 2018). Some systems integrate float switches or conductive tilt detection to identify bottle usage (Pankajavalli et al., 2017; Ko et al., 2007). Ambient humidity has also been used to characterize dining environments and inform food safety assessments (Bastida et al., 2023; Badawi et al., 2012b; Kwon et al., 2016).
@@
Magnetic technologies enable automatic identification of food items, storage states, or utensil presence (Vazquez et al., 2015; Huang et al., 2010; Jung et al., 2017). Some advanced RFID or NFC designs estimate consumed volume via passive milliliter tracking (Yıldız and Coşkun, 2019). Additionally, Han and Kang (2017) used metal contact detection to support real-time eating activity sensing.
5.2. Intervention